Solving Quadratic Equations by Completing the Square
1.x^2+6x=16
2.x^2+4x=45
3.x^2+8x=-8
4.x^2-12x=-18
5.x^2-10x=50

Answer by mananth(16946)

x^2+6x=16
Finding the third term to complete the square
Third term = ( (1/2) * co-efficient of x)^2
= (( 1/2) *6)^2
=9
x^2+6x=16
add 9 to both sides
x^2+6x+9=16+9
(x+3)^2 = 25
x+3 = +/- 5
x= -3+5 = +2
OR
x=-3-5 =-8
x= 2 OR -8